PC Planet has just opened its doors. The new retail store sells refurbished computers at a significant discount from market prices. The computers cost PC Planet $100 to purchase and require 10 hours of labor at $15 per hour. Additional variable costs, including wages for sales personnel, are $50 per computer. The newly refurbished computers are resold to customers for $500. Rent on the retail store costs the company $4,000 per month.

1. How many computers does PC Planet have to sell each month to break even?
2. If PC Planet wants to earn $5,000 per month after all expenses, how many computers does the company need to sell?
3. PC Planet can purchase already refurbished computers for $200. This would mean that all labor required to refurbish the computers could be eliminated. What would PC Planet’s new breakeven point be if it decided to purchase the computers already refurbished?
4. Instead of paying the monthly rental fee for the retail space, PC Planet has the option of paying its landlord a 20% commission on sales. Assuming the original facts in the problem, at what sales level would PC Planet be indifferent between paying a fixed amount of monthly rent and paying a 20% commission on sales?
Answer #1
Contribution margin per computer = Selling price per Computer - variable costs per computer
=500-100-10*15-50
200 per computer
Number of computers to break even = Fixed costs/Contribution margin per computer
=4000/200
20 computers
2.Number of computers = (Desired income + fixed costs)/Contribution margin per computer
=(4000+5000)/200
45 computers
3.Break even point = 4000/(500-200-50)
=16 computers
4.Let the sales level be x
20%*500*x = 4000
x = 40 computers